OpenStudy (anonymous):

Is this equation linear in the variables of x and y: (sin2)x-y=14

OpenStudy (anonymous):

My answer was yes, anyone please let me know if this is right

jimthompson5910 (jim_thompson5910):

if the equation is \[\Large \sin(2)*x - y = 14\] then you are correct the sin(2) part is a number, approx 0.034899 so the equation is approximately equivalent to 0.034899x - y = 14. You can place it into y = mx+b form

OpenStudy (anonymous):

Okay thank you, I also had another question. Why is 2sinx-y=14 non linear?

jimthompson5910 (jim_thompson5910):

because of the sin(x) term if you solve for y, you get y = 2sin(x) - 14 which graphs a curvy/wavy like equation and NOT a straight line. https://www.desmos.com/calculator/6xg8ycssnj

OpenStudy (anonymous):

oh okay got it, thank you

jimthompson5910 (jim_thompson5910):

the big difference is what goes inside the sine function if it's a single number that doesn't change, then you can replace the sine expression with some fixed number that doesn't change but if x or some algebraic expression is inside the sine function, then you'll get that curve and pattern

jimthompson5910 (jim_thompson5910):

no problem